Which of the following is not considered an automatic stabilizer?

A.

food stamp program for people with low incomes

B.

welfare program for families with dependent children

C.

Medicaid, a health program for the poor

D.

financial assistance for disabled people

E.

unemployment programs that pay benefits to those who lose their jobs

If the MPC is 0.75, then the spending multiplier is:

A.

-o.57

B.

0.57

C.

4.0

D.

-4.0

Which of the following Federal Reserve policy tools is used most often?

A.

open market operations

B.

changing the required reserve ratio

C.

changing the discount rate

D.

changing the prime lending rate

E.

moral suasion

Automatic stabilizers are:

A.

well-timed government policy changes that help keep the economy stable at full employment.

B.

forces in the economy that increase the size of the expenditure multiplier.

C.

spending shocks that move in the opposite direction of, and help to neutralize, other spending shocks.

D.

actions taken by foreign governments that help stabilize our own economy.

E.

All of these

F.

None of these
Answer #1

1. Option D.

  • Automatic stabilizers are those tools that function like the fiscal policy but function without any government intervention to stabilize the economy during any Economic down turns.
  • Examples of automatic stabilizers include food stamps, Medicaid programs, unemployment benefits, welfare programs etc.
  • Financial assistance For disabled people is not considered as an automatic stabilizer.

2. Option C.

  • Given MPC = 0.75, then spending multiplier = 1/1-MPC.
  • Spending multiplier = 1/1-0.75 = 4//

3. Option A.

  • Open market operations are the most commonly used monetary policy tool by Fed.
  • Through open market operations the fed can either Increase or decrease the money supply within the economy.

4. Option F.

  • None of the above options are the definitions of automatic stabilizers.
  • Automatic stabilizers are the tools that stabilize an economy automatically without any government intervention.